# A Closed-form Solution to the Wahba Problem for Pairwise Similar Quaternions

### v1

@@ Abstract.
In this work, we derive an explicit analytical characterization of all quaternions that yield zero Wahba cost for the case $\ell=2$. Our approach builds on a connection between quaternion similarity, the singular Sylvester equation aq=qb{a}{q}={q}{b}, and quaternion square roots established in our previous work [1]. We provide (i) necessary and sufficient conditions under which the Wahba’s cost function is zero and (ii) a closed-form parameterization of all such quaternions. This eliminates the need for eigenvalue computations and enables a direct algebraic understanding of the underlying geometry of Wahba’s problem.
@@ 1. Introduction
Determining the relative orientation between two Cartesian coordinate systems is a fundamental in many practical applications, such as spacecraft attitude estimation, robotics, and autonomous driving. These applications require fast and accurate information about the object’s orientation. Solutions often rely on vector observations expressed in two coordinate systems. The problem of finding the relationship between two sets of vector measurements is formally known as the Wahba’s problem [2]. Wahba’s problem has been addressed using various mathematical formulations, including rotation matrices, direction cosine matrices, and quaternions. Among these, quaternion-based approaches are particularly effective due to their compact representation, absence of singularities, and computational efficiency.
Wahba’s problem in quaternions states: Given two sets of pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{\ell}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{\ell}\}$, where $\ell\geq 2$, Wahba’s problem aims at finding a nonzero quaternion q{q} that minimizes the cost function
@@
Pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{\ell}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{\ell}\}$ represent vectors observed in the sensor frames $\mathcal{\!}{A}$ and $\mathcal{\!}{B}$, respectively.
Several well-established algorithms solve Wahba’s problem, including Davenport’s q{q}-method [3], the QUEST algorithms [4, 5, 6], Horn’s formulation [7], and ESOQ methods [8, 9] for the general case $\ell\geq 2$. All of these methods obtain the optimal quaternion either through the eigendecomposition of a $4\times 4$ matrix or through iterative numerical procedures. As a result, the solution is obtained indirectly via numerical linear algebra, without exposing the analytic relationship between the structure of Wahba’s cost function and the form of its optimal quaternion. Consequently, to the best of our knowledge, the literature contains no closed-form quaternion solution that fully exploits the special structure of the case $\ell=2$.
In this paper, we solve the Wahba’s problem for the special and practically important case $\ell=2$. We establish a direct relation between the quaternionic linear equation aqqb=0{a}{q}-{q}{b}=0, commonly known as the Sylvester equation, studied in our previous work [1], and the Wahba’s cost function in (1.1). This relation leads to the two main contributions of this work: (i) the necessary and sufficient condition for the existence of solutions and (ii) an analytical expression for the solution set of f(q)=0f({q})=0 when $\ell=2$. Our approach avoids conventional matrix-based techniques, such as iterative eigenvalue computation or matrix inversion, thereby improving computational efficiency and analytical clarity of the solution.
The remainder of the paper is structured as follows. Section 2 reviews the necessary quaternionic preliminaries, introduces a notion of pairwise similarity, and summarizes background results from [1]. Section 4 presents the main results: the existence criterion and the complete closed-form solution set of f(q)=0f({q})=0 when $\ell=2$. Section 5 concludes the paper.
@@ 2. Preliminaries
This section introduces the quaternion notation and reviews the relevant algebraic properties of quaternions used throughout the paper. In particular, the notion of pairwise quaternion similarity is introduced, which will be the central concept governing the solvability of Wahba’s cost function for $\ell=2$.
@@ 4. Solutions to Wahba’s problem in quaternions
In this section, we establish the connection between the solvability of (4.1) and quaternion pairwise similarity. We show that (4.1) corresponds to pairs of singular Sylvester equations, which underpins the algebraic structure of the Wahba’s problem for the special case $\ell=2$. Based on this, we show the necessary and sufficient conditions for existance of solution to (4.1) and derive their explicit analytical representation.
@@
Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions with $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$ and $\lambda_{1},\lambda_{2}$, and $\mu$ real numbers with $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}\,\mathrm{Im}\,{({a}_{1}+{b}_{1})}\mathclose{\rvert}\neq 0$ and $\lambda_{2}\neq 0$. Then, the general solution to (4.1) is given as ${q}={q}_{1}{q}_{2}$, where
@@
Let ${q}_{1}$ be a nonzero quaternion such that $\mathopen{\lvert}{q}_{1}^{-1}{a}_{1}{q}_{1}-{b}_{1}\mathclose{\rvert}^{2}=0$. Then, it equivalently holds ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$. Moreover, by Corollary 4.1.1 we obtain ${q}_{1}$ in (4.4) subject to $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu\,({a}_{1}+{b}_{1})\mathclose{\rvert}\neq 0$ and ${a}_{1}\cdot{\mathrm{Im}\,{q}}_{1}=0$ if ${a}_{1}=-{b}_{1}$. For any nonzero quaternion q{q} there exists a unique nonzero quaternion ${q}_{2}$ such that ${q}={q}_{1}{q}_{2}$. Then, substituting ${q}={q}_{1}{q}_{2}$ and ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$ into the left side equality in (4.6) yields
@@
Define ${a}_{3}={q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}$ and ${b}_{3}={b}_{1}\times{b}_{2}$. Then, by the left side property in (2.1), and since $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$, we obtain
@@ 5. Conclusion
This work establishes the necessary and sufficient condition under which Wahba’s cost function for two vector observations equals zero. Subsequently, we derived the corresponding closed-form solution for pairwise similar quaternions, by revealing the connection between the singular Sylvester equation in quaternions and quaternionic square roots. We obtained a direct analytical pathway to the closed-form solution in the two-observation case. The presented results offer a complete theoretical characterisation of the Wahba problem in quaternions and lay a rigorous foundation for further generalisations in quaternion-based orientation estimation.

### v2

# A Closed-form Solution to the Wahba’s Problem for Pairwise Similar Quaternions

@@ Abstract.
In this work, we derive an explicit analytical characterization of all quaternions that yield zero Wahba’s cost for the case $n=2$. Our approach builds on a connection between quaternion similarity, the singular Sylvester equation aq=qb{a}{q}={q}{b}, and quaternion square roots established in our previous work [5]. We provide (i) necessary and sufficient conditions under which the Wahba’s cost function is zero and (ii) a closed-form parameterization of all such quaternions. This eliminates the need for eigenvalue computations and enables a direct algebraic understanding of the underlying geometry of Wahba’s problem.
Keywords: Quaternions; Wahba’s problem; Similarity; Pairwise similarity; Closed-form solution
Mathematical sub classification numbers (MSC) 2020: 15A24, 15B33
@@ 1. Introduction
Determining the relative orientation between two Cartesian coordinate systems is a fundamental in many practical applications, such as spacecraft attitude estimation, robotics, and autonomous driving. These applications require fast and accurate information about the object’s orientation. Solutions often rely on vector observations expressed in two coordinate systems. The problem of finding the relationship between two sets of vector measurements is formally known as Wahba’s problem [10]. Wahba’s problem has been addressed using various mathematical formulations, including rotation matrices, direction cosine matrices, and quaternions. Among these, quaternion-based approaches are particularly effective due to their compact representation, absence of singularities, and computational efficiency.
Wahba’s problem in quaternions states: Given two sets of pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{n}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{n}\}$, where $n\geq 2$, Wahba’s problem aims at finding a nonzero quaternion q{q} that minimizes the cost function
@@
Pure quaternions $\{{a}_{1},{a}_{2},\dots,{a}_{n}\}$ and $\{{b}_{1},{b}_{2},\dots,{b}_{n}\}$ represent vectors observed in the sensor frames $\mathcal{\!}{A}$ and $\mathcal{\!}{B}$, respectively.
Several well-established algorithms solve Wahba’s problem, including Davenport’s q{q}-method [1], the QUEST algorithms [8, 9, 7], Horn’s formulation [2], and ESOQ methods [4, 3] for the general case $n\geq 2$. All of these methods obtain the optimal quaternion either through the eigendecomposition of a $4\times 4$ matrix or through iterative numerical procedures. As a result, the solution is obtained indirectly via numerical linear algebra, without exposing the analytic relationship between the structure of Wahba’s cost function and the form of its optimal quaternion. Consequently, to the best of our knowledge, the literature contains no closed-form quaternion solution that fully exploits the special structure of the case $n=2$.
In this paper, we solve the Wahba’s problem for the special and practically important case $n=2$. We establish a direct relation between the quaternionic linear equation aqqb=0{a}{q}-{q}{b}=0, commonly known as the Sylvester equation, studied in our previous work [5], and the Wahba’s cost function in (1.1). This relation leads to the two main contributions of this work: (i) the necessary and sufficient condition for the existence of solutions and (ii) an analytical expression for the solution set of f(q)=0f({q})=0 when $n=2$. Our approach avoids conventional matrix-based techniques, such as iterative eigenvalue computation or matrix inversion, thereby improving computational efficiency and analytical clarity of the solution.
The remainder of the paper is structured as follows. Section 2 reviews the necessary quaternionic preliminaries, introduces a notion of pairwise similarity, and summarizes background results from [5]. Section 4 presents the main results: the existence criterion and the complete closed-form solution set of f(q)=0f({q})=0 when $n=2$. Section 5 concludes the paper.
@@ 2. Preliminaries
This section introduces the quaternion notation and reviews the relevant algebraic properties of quaternions used throughout the paper. In particular, the notion of pairwise quaternion similarity is introduced, which will be the central concept governing the solvability of Wahba’s cost function for $n=2$.
@@ 4. Solutions to Wahba’s problem in quaternions
In this section, we establish the connection between the solvability of (4.1) and quaternion pairwise similarity. We show that (4.1) corresponds to pairs of singular Sylvester equations, which underpins the algebraic structure of the Wahba’s problem for the special case $n=2$. Based on this, we show the necessary and sufficient conditions for existance of solution to (4.1) and derive their explicit analytical representation.
@@
Let ${a}_{1},{a}_{2},{b}_{1},{b}_{2}$ be nonreal quaternions with $({a}_{1},{a}_{2})\sim({b}_{1},{b}_{2})$ and $\lambda_{1},\lambda_{2}$, and $\mu_{1}$ real numbers with $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}\,\mathrm{Im}\,{({a}_{1}+{b}_{1})}\mathclose{\rvert}\neq 0$ and $\lambda_{2}\neq 0$. Then, the general solution to (4.1) is given as ${q}={q}_{1}{q}_{2}$, where
@@
Let ${q}_{1}$ be a nonzero quaternion such that $\mathopen{\lvert}{q}_{1}^{-1}{a}_{1}{q}_{1}-{b}_{1}\mathclose{\rvert}^{2}=0$. Then, it equivalently holds ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$. Moreover, by Corollary 4.1.1 we obtain ${q}_{1}$ in (4.4) subject to $\mathopen{\lvert}\lambda_{1}\mathclose{\rvert}+\mathopen{\lvert}\mu_{1}({a}_{1}+{b}_{1})\mathclose{\rvert}\neq 0$ and ${a}_{1}\cdot{\mathrm{Im}\,{q}}_{1}=0$ if ${a}_{1}=-{b}_{1}$. For any nonzero quaternion q{q} there exists a unique nonzero quaternion ${q}_{2}$ such that ${q}={q}_{1}{q}_{2}$. Then, substituting ${q}={q}_{1}{q}_{2}$ and ${a}_{1}{q}_{1}={q}_{1}{b}_{1}$ into the left side equality in (4.6) yields
@@
Define ${a}_{3}={q}_{1}^{-1}({a}_{1}\times{a}_{2}){q}_{1}$ and ${b}_{3}={b}_{1}\times{b}_{2}$. Then, by the left side property in (2.1) and since $\mathrm{Re}\,({a}_{1}{a}_{2})=\mathrm{Re}\,({b}_{1}{b}_{2})$, we obtain
@@ 5. Conclusion
This work establishes the necessary and sufficient condition under which Wahba’s cost function for two vector observations equals zero. Subsequently, we derived the corresponding closed-form solution for pairwise similar quaternions by revealing the connection between the singular Sylvester equation in quaternions and quaternionic square roots. We obtained a direct analytical pathway to the closed-form solution in the two-observation case. The presented results offer a complete theoretical characterisation of the Wahba’s problem in quaternions and lay a rigorous foundation for further generalisations in quaternion-based orientation estimation.